OpenStudy (anonymous):

How do you simplify 1/(cos(x/2)sin(x/2)) to 2cosec(x) ?

OpenStudy (stacey):

Use \[\sin(2\theta)=2\sin \theta \cos \theta\]

OpenStudy (stacey):

\[\theta=x/2\]

OpenStudy (anonymous):

\[\frac{1}{\text{Cos}\left[\frac{x}{2}\right]\text{Sin}\left[\frac{x}{2}\right]}=\frac{1}{\left(\frac{\text{Sin}[x]}{2}\right)}=2 \text{Csc}[x] \]

OpenStudy (anonymous):

\[\frac{1}{\text{Cos}\left[\frac{x}{2}\right]\text{Sin}\left[\frac{x}{2}\right]}=\frac{1}{\left(\frac{\text{Sin}[x]}{2}\right)}=2 \text{Csc}[x] \]